Suppose a new MyPhone cellphone costs $650, and depreciates in value by $30 per month. A new Cyborg phone costs $580, and depreciates in value by $25 per month. 


 (a) Let M(x) be the value of a MyPhone r months after it was purchased, and let C(x) be the value of a Cyborg a months after it was purchased. Write formulas for M(x) and C(x). 


 (b) If Carrie and Miranda buy phones at the same time, Carrie buying the Cyborg and Miranda buying the MyPhone, when will their phones have the same value?


 (c) After how many months are each of the phones worth nothing?
Answer #1

W M - 650 - 30x CO) - 580-25% 6) M(x) =C(x) 650 -30%= 580-25% 70= 5x * x = 14 months c M(x) = 0 = 650-30x=0 > x= 650 - 21-67

\blacksquare

Hope this was helpful. Please do leave a positive rating if you liked this answer. Thanks and have a good day!